STATUS ON THE GLOBAL VIBRATION-ROTATION MODEL IN ACETYLENE

B. AMYAY, \underline{M. HERMAN}, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium; A. FAYT, Laboratoire de Spectroscopie Moléculaire, Université Catholique de Louvain, Chemin du Cyclotron, 2, B-1348 Louvain-La-Neuve, Belgium.

We have developed a global model to deal with all vibration-rotation levels in acetylene up to high vibrational excitation energy, typically up to 9000 wavenumbers. It has been applied to a number of isotopologues, considering all known vibration-rotation lines published in the literature, for various purposes such as line assignment^{a} and astrophysical applications^{b}^{b}. Coriolis interaction is now systematically being introduced in the model. Recent results concerning the analysis of hot emission FTIR spectra recorded around 3 microns by R. Georges et al. at the University of Rennes (France) and of CW-CRDS spectra recorded around 1.5 microns by A. Campargue et al. at the University of Grenoble (France) will help illustrate the role of this vibration-rotation coupling in the global polyad scheme.

aS. Robert, M. Herman, A. Fayt, A. Campargue, S. Kassi, A. Liu, L. Wang, G. Di Lonardo, and L. Fusina, Mol. Phys., \underline{106}, 2581 (2008).

bA. Jolly, Y. Benilan, E. Cané, L. Fusina, F. Tamassia, A. Fayt, S. Robert, and M. Herman, J.Q.S.R.T., \underline{109}, 2846 (2008).